A solution is prepared by mixing 1 g of ethanol (C2H6O) with 100 g water to give a final

volume of 101 ml. Calculate the mole fraction for the solute and solvent, and the molality of

ethanol in the solution.


1
Expert's answer
2022-02-23T23:51:01-0500

Molar mass of ethanol=46

Molar mass of water =18

Moles of ethanol=1/46= 0.02174

Moles of water=100/18=5.55555

Mole fraction of C2H6O= 0.02174/(0.02174+5.55555)

=0.02174/5.57729

=0.004

Mole fraction of water =5.55555/5.57729

=0.996


Molality of ethanol=(1/46)/(100/1000)

=0.02174/0.1

=0.2174